OpenStudy (anonymous):

a sphere of mass 40 kg is atracted by a second sphere of mass15 kg when their centres are 20 cm apart,with a force of 0.1milligram weight. calculate the value of gravitational constant.

OpenStudy (anonymous):

plzz doo

OpenStudy (hba):

What is Newton's Universal Law of Gravitation? That is what you have to use for this problem.

OpenStudy (anonymous):

i dont knw...

OpenStudy (hba):

gravitational force = 0.1 mg-wt. = 0.1 * 10^-6 * 9.8 N = Gm1m2/r^2 m1 = 40 kg m2 =15 kg and r = 0.2 m Put in and find G

OpenStudy (hba):

F=Gm1m2/r^2

OpenStudy (anonymous):

kya ap mujhe samjha sakte ho???

OpenStudy (hba):

what ?

OpenStudy (anonymous):

the ans that u hv written

OpenStudy (hba):